# Topological symmetry groups of the Generalized Petersen Graphs

### v1

@@ Abstract.
The topological symmetry group $\mathrm{TSG}(\Gamma)$ of an embedding $\Gamma$ of a graph in $S^{3}$ is the subgroup of the automorphism group of the graph which is induced by homeomorphisms of $(S^{3},\Gamma)$. If we restrict to orientation preserving homeomorphisms then we obtain the orientation preserving topological symmetry group $\mathrm{TSG}_{+}(\Gamma)$. In this paper we determine all groups that can be $\mathrm{TSG}(\Gamma)$ or $\mathrm{TSG}_{+}(\Gamma)$ for some embedding $\Gamma$ of a generalized Petersen graph other than the exceptional graphs $P(12,5)$ and $P(24,5)$ (which will be addressed in a separate paper.
@@ 1. Introduction
The field of spatial graph theory developed in part to classify the symmetries of flexible molecules [14] and in part as an extension of knot theory [4]. Knot theorists have a long tradition of classifying the symmetries of knots and links. However, in contrast with the symmetries of a knot or link, the symmetries of a spatial graph can be understood in terms of the automorphism group of its underlying graph. In particular, we have the following definitions.
@@
Various results have been obtained about the realizable and positively realizable groups for specific graphs (see for example [3],[6],[7],[8]). In this paper, we classify the groups that are realizable and positively realizable for the family of generalized Petersen graphs. In particular, for $2k<n$, the generalized Petersen graph $P(n,k)$ is obtained from an $n$-gon with consecutive vertices $u_{1},u_{2},\dots,u_{n}$ and a star (possibly with more than one loop) with vertices $v_{1},v_{2},\dots,v_{n}$ and edges $\overline{v_{i}v_{i+k}}$, by adding an edge $\overline{u_{i}v_{i}}$ for each $i\leq n$ (see for example, the graph $P(6,2)$ and an embedding of $P(7,2)$ in Figure 1). We call the edges $\overline{u_{i}u_{i+1}}$ the outer edges, the edges $\overline{v_{i}v_{i+k}}$ the inner edges, and the edges $\overline{u_{i}v_{i}}$ the spokes. Note that when $k^{2}\equiv\pm 1\pmod{n}$, then $k$ and $n$ are relatively prime and hence the star formed by the inner edges of $P(n,k)$ is a single cycle.
@@
Using the above, we classify the realizable and positively realizable subgroups of $\mathrm{Aut}(P(n,k))$ for all non-exceptional pairs $(n,k)$ and for the exceptional pairs $(4,1)$, $(8,3)$, $(10,2)$, $(10,3)$. For the exceptional pair $(5,2)$, the graph $P(5,2)$ is the Petersen graph whose topological symmetry groups were determined by Chambers, Flapan, Heath, Lawrence, Thatcher, and Vanderpool [2]. In particular, they showed that not all subgroups of $\mathrm{Aut}(P(5,2))$ are realizable or positively realizable. Realizability and positive realizability for $(12,5)$ and $(24,5)$ will be considered in a subsequent paper.
Observe that if $G=\mathrm{TSG}_{+}(\Gamma)$ for some embedding $\Gamma$ of $P(n,k)$, then we can add the same chiral invertible knot to every edge of $\Gamma$ to get an embedding $\Gamma^{\prime}$ with $G=\mathrm{TSG}(\Gamma^{\prime})$. This means that if we can find an embedding of $P(n,k)$ such that $G=\mathrm{TSG}_{+}(\Gamma)$, then $G$ is both positively realizable and realizable for $P(n,k)$. Also, by putting a distinct knot on each edge of $\Gamma$ we obtain an embedding $\Gamma^{\prime\prime}$ such that $\mathrm{TSG}(\Gamma^{\prime\prime})=\mathrm{TSG}_{+}(\Gamma^{\prime\prime})=\langle e\rangle$. Thus the trivial group is realizable and positively realizable for every $P(n,k)$. Since $P(n,k)$ is $3$-connected, we can use the following theorems.
@@
The theorem below follows from Proposition 3 of [9] together with the Geometrization Theorem [13]. Note that $\mathrm{SO}(4)$ is the group of orientation preserving isometries of $S^{3}$ and $\mathrm{SO}(3)$ is the group of orientation preserving isometries of $S^{2}$.
@@
Let $\gamma$ be a $3$-connected graph such that $G\leq\mathrm{TSG}_{+}(\Lambda)$ for some embedding $\Lambda$ of $\gamma$ in $S^{3}$. Then there is an embedding $\Gamma$ of $\gamma$ in $S^{3}$ such that $G\leq\mathrm{TSG}_{+}(\Gamma)$ and $\mathrm{TSG}_{+}(\Gamma)$ is induced by an isomorphic finite subgroup of $\mathrm{SO}(4)$.
@@
Let $G\leq SO(4)$ such that for every involution $g\in G$, we have $\mathrm{fix}(g)\cong S^{1}$ and no $h\in G$ which is distinct from $g$ has $\mathrm{fix}(h)=\mathrm{fix}(g)$. Then $G$ is either a subgroup of $D_{m}\times D_{m}$ for some odd $m$ or a finite subgroup of $\mathrm{SO}(3)$.
@@
Then there is an embedding of the edges of $\gamma$ in $S^{3}$ such that the resulting embedding of $\gamma$ is setwise invariant under $G$.
We also use the results of Frucht et al. [12], the software Sage, and the website, https://people.maths.bris.ac.uk/ matyd/GroupNames/ hosted by the University of Bristol to identify the isomorphism classes of automorphism groups of the generalized Petersen graphs and their subgroups.
2. Realizability of $D_{n}$ and its subgroups
@@
We construct an embedding of $P(n,k)$ as follows. Let $0<r<R$. We begin with a planar circle $C$ of radius $R$ centered at the origin. Then embed each $u_{i}$ at the intersection of $C$ with a ray at an angle of $\frac{\pi}{2}+\frac{2\pi}{n}(i-1)$, and embed each $v_{i}$ at the intersection of a planar circle $c$ of radius $r$ centered at the origin with a ray at an angle of $\frac{\pi}{2}+\frac{2\pi}{n}(i-1)$. Then we embed the spokes $\overline{u_{i}v_{i}}$ and outer edges $\overline{u_{i}u_{i+1}}$ as straight line segments.
For each $i$, we add a straight segment between $v_{i}$ and $v_{i+k}$. If $k=1$, then there are no crossings between these segments, and hence we let these segments be the inner edges $\overline{v_{i}v_{i+k}}$ which together form an $n$-gon. If $k>1$, then this segment crosses $22k-$ 2 other inner edges. In this case, we embed the inner edge $\overline{v_{i}v_{i+k}}$ so that its shadow is this segment and its crossings with the other inner edges alternate, beginning with an overcrossing closest to $v_{i}$. Let $\Gamma^{\prime}$ denote this embedding of $P(n,k)$ (see the embedding of $P(7,2)$ on the right of Figure 1). We will construct our embedding $\Gamma$ from $\Gamma^{\prime}$ as described below.
Let $f$ denote a rotation of $S^{3}$ of $\frac{2\pi}{n}$ about an axis through the origin which is perpendicular to the plane of the projection, and let $g$ denote a rotation of $\pi$ about a planar axis that contains the spoke $\overline{u_{n}v_{n}}$ if $n$ is odd and contains the two opposite spokes $\overline{u_{n}v_{n}}$ and $\overline{u_{\frac{n}{2}}v_{\frac{n}{2}}}$ if $n$ is even. Then the group $\langle f,g\rangle$ leaves the circle of outer edges and the set of spokes invariant, inducing the dihedral group $D_{n}$ on the vertices of $\Gamma^{\prime}$.
If $k=1$, then $\langle f,g\rangle$ also leaves the $n$-gon of inner edges invariant. Suppose that $k>1$. Then because each $\overline{v_{i}v_{i+k}}$ has the same pattern of crossings with the other inner edges, the star made up of inner edges is invariant under the rotation $f$. Also, $\beta(v_{i})\equiv v_{-i}\pmod{n}$, and hence $g$ maps the edge $\overline{v_{i}v_{i+k}}$ to the edge $\overline{v_{-i}v_{-i-k}}$ with the appropriate crossings. It follows that the set of inner edges is also preserved under $g$. Thus $\Gamma^{\prime}$ is invariant under $\langle f,g\rangle$, and hence $D_{n}\leq\mathrm{TSG}_{+}(\Gamma^{\prime})$.
We obtain the embedding $\Gamma$ from $\Gamma^{\prime}$ by adding the invertible knot $4_{1}$ to each outer edge. Then $\Gamma$ is still invariant under $\langle f,g\rangle$, and hence $D_{n}\leq\mathrm{TSG}_{+}(\Gamma)$. Let $L$ denote the $n$-gon of outer edges of $\Gamma^{\prime}$. Then $L$ is the only $n$-gon containing $n$ copies of the knot $4_{1}$, and hence any element of $\mathrm{TSG}_{+}(\Gamma)$ must take $L$ to itself. Since no non-trivial automorphism of $P(n,k)$ fixes every $u_{i}$, this implies that $\mathrm{TSG}_{+}(\Gamma)\leq\mathrm{TSG}_{+}(L)$. On the other hand, because the automorphism group of an $n$-gon is $D_{n}$, we have $\mathrm{TSG}_{+}(L)\leq D_{n}$, and hence $D_{n}=\mathrm{TSG}_{+}(\Gamma)$.
@@
Let $k^{2}\not\equiv\pm 1\pmod{n}$ where $(n,k)$ is non-exceptional. Then $\operatorname{Aut}(P(n,k))=D_{n}$ and all of its subgroups are positively realizable and hence realizable for $P(n,k)$.
@@ 3. The case $k^{2}\equiv 1\pmod{n}$
We will use the following embedding $\Lambda$ of $P(n,k)$ in this and the next section. Let $U$ and $V$ denote the cores of complementary isometric solid tori in $S^{3}$. Hence $U$ and $V$ are geodesic circles, and for every point on one of these cores, its antipodal point is on the same core. Thus for every $u\in U$ and $v\in V$, there is a unique shortest geodesic $e$ joining $u$ and $v$, and the length of $e$ is less than $\pi$. It follows that the only shortest geodesic between a pair of points on $e$ is contained in $e$. In particular, since $U$ and $V$ are geodesic circles, the interior of $e$ must be disjoint from $U\cup V$.
Suppose that $k^{2}\equiv\pm 1\pmod{n}$. Then $n$ and $k$ are relatively prime and the inner edges of $P(n,k)$ form a single loop. The embedding $\Lambda$ is obtained by letting $U$ and $V$ be the outer cycle and inner cycle, respectively, of $\Lambda$ and putting $n$ evenly spaced vertices on each of $U$ and $V$, labeling them so that edges on $U$ have the form $\overline{u_{i}u_{i+1}}$ and edges on $V$ have the form $\overline{v_{i}v_{i+k}}$. Then for each $i$, we let $e_{i}$ be the unique shortest geodesic between $u_{i}$ and $v_{i}$. Orient $U$ and $V$ so that the edges $\overline{u_{i}u_{i+1}}$ and $\overline{v_{i}v_{i+k}}$ are oriented positively.
@@
We start with the embedding $\Lambda$ described above and we add a $4_{1}$ knot to each edge of $U$ and $V$ to get an embedding $\Lambda^{\prime}$ of $P(n,k)$. Let $f$ be a glide rotation which rotates $U$ by $\frac{2\pi}{n}$ while rotating $V$ by $\frac{2k\pi}{n}$ both in the positive direction. Let $g$ be a rotation of $S^{3}$ by $\pi$ around an axis that contains $e_{1}=\overline{u_{1}v_{1}}$. Then $g$ takes $U$ and $V$ to themselves, reversing their orientations. Finally, let $h$ be an order $2$ rotation of $S^{3}$ interchanging the positively oriented $U$ with the positively oriented $V$, taking $u_{n}$ to $v_{n}$.
In order to show that $f$, $g$, and $h$ take edges to edges, first observe that $f(u_{i})=u_{i+1}$ and $f(v_{i})=v_{i+k^{2}}=v_{i+1}$. Since $e_{i+1}$ is the unique shortest geodesic between $u_{i+1}$ and $v_{i+1}$, it follows that $f(e_{i})=e_{i+1}$. Also, for every $i$, we have $g(v_{1-ki})=v_{1+ki}$, and hence $g(v_{1-i})=v_{1+i}$. Since we also have $g(u_{1-i})=u_{1+i}$, it follows that $g(e_{1-i})=e_{i+1}$. Finally, for each $i$, we have $h(u_{i})=v_{ki}$, and hence $h(u_{ki})=v_{k^{2}i}=v_{i}$. Since $h$ has order $2$, it follows that $h(v_{i})=u_{ki}$. Thus $h(e_{i})=e_{ki}$. Thus $f$, $g$, and $h$ take $\Lambda^{\prime}$ to itself.
Observe that $f$ induces a rotation of order $n$ on $U$ and $V$, $g$ turns both $U$ and $V$ over, and $h$ is an involution interchanging $U$ and $V$. Thus the group that $\langle f,g,h\rangle$ induces on $\Lambda^{\prime}$ is in fact $D_{n}\rtimes\mathbb{Z}_{2}$, and hence $D_{n}\rtimes\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Lambda)$. However, since every element of $\mathrm{TSG}_{+}(\Lambda)$ takes the set of edges $e_{i}$ to themselves, it follows that $\mathrm{TSG}_{+}(\Lambda^{\prime})\leq B(n,k)$. Now by [12], $B(n,k)=D_{n}\rtimes\mathbb{Z}_{2}$ and hence $\mathrm{TSG}_{+}(\Lambda^{\prime})=D_{n}\rtimes\mathbb{Z}_{2}$.
@@
Since we know by [12] that for $(n,k)$ non-exceptional, $\mathrm{Aut}(P(n,k))=B(n,k)$, we have the following.

### v2

@@ Abstract.
The topological symmetry group $\mathrm{TSG}(\Gamma)$ of an embedded graph $\Gamma$ in $S^{3}$ is the subgroup of the automorphism group of the graph which is induced by homeomorphisms of $(S^{3},\Gamma)$. If we restrict to orientation preserving homeomorphisms then we obtain the orientation preserving topological symmetry group $\mathrm{TSG}_{+}(\Gamma)$. In this paper, we determine all groups that can be $\mathrm{TSG}(\Gamma)$ or $\mathrm{TSG}_{+}(\Gamma)$ for some embedding $\Gamma$ of a generalized Petersen graph other than the exceptional graphs $P(12,5)$ and $P(24,5)$.
@@ 1. Introduction
The field of spatial graph theory was developed in part to classify the symmetries of flexible molecules [14] and in part as an extension of the study of the symmetries of knots and links [4]. However, in contrast with the symmetries of a knot or link, the symmetries of a spatial graph can be understood in terms of the automorphism group of its underlying graph. In particular, we have the following definitions.
@@
Many previous results have been obtained about realizable and positively realizable groups for specific graphs (see for example [3],[6],[7],[8]). In this paper, we classify the groups that are realizable and positively realizable for the family of generalized Petersen graphs. In particular, for $2k<n$, the generalized Petersen graph $P(n,k)$ is obtained from an $n$-gon with consecutive vertices $u_{1},u_{2},\dots,u_{n}$ and a star (possibly with more than one loop) with vertices $v_{1},v_{2},\dots,v_{n}$ and edges $\overline{v_{i}v_{i+k}}$, by adding an edge $\overline{u_{i}v_{i}}$ for each $i\leq n$ (see the graph $P(6,2)$ and an embedding of $P(7,2)$ in Figure 1). We call the edges $\overline{u_{i}u_{i+1}}$ outer edges, the edges $\overline{v_{i}v_{i+k}}$ inner edges, and the edges $\overline{u_{i}v_{i}}$ spokes. Note that when $k^{2}\equiv\pm 1\pmod{n}$, then $k$ and $n$ are relatively prime and hence the star formed by the inner edges of $P(n,k)$ is a single cycle.
@@
We classify the realizable and positively realizable subgroups of $\mathrm{Aut}(P(n,k))$ for all non-exceptional pairs $(n,k)$ and for the exceptional pairs $(4,1)$, $(8,3)$, $(10,2)$, $(10,3)$. For the exceptional pair $(5,2)$, the graph $P(5,2)$ is the Petersen graph whose topological symmetry groups were determined by Chambers, Flapan, Heath, Lawrence, Thatcher, and Vanderpool [2]. Realizability and positive realizability for $P(12,5)$ and $P(24,5)$ will be considered in a subsequent paper. The exceptional graphs $P(n,k)$ for $n\leq 10$ are illustrated in Figure 2.
Figure 2. The exceptional graphs $P(4,1)$, $P(5,2)$, $P(8,3)$, $P(10,2)$, and $P(10,3)$.
Our arguments make use of the above theorem of Frucht et al, as well as the theorems listed below. However, for a large number of groups, we prove realizability or positive realizability by constructing intricate embeddings which we show have the required group of symmetries and not a larger group (the latter is often the difficult part). These embeddings are described in the text, and augmented by detailed illustrations. Some readers will find these constructions to be the most interesting part of the paper, since they can likely be generalized to other graphs.
If $G=\mathrm{TSG}_{+}(\Gamma)$ for some embedding $\Gamma$ of $P(n,k)$, then we can add the same chiral invertible knot to every edge of $\Gamma$ to get an embedding $\Gamma^{\prime}$ with $G=\mathrm{TSG}(\Gamma^{\prime})$. This means that if we can find an embedding of $P(n,k)$ such that $G=\mathrm{TSG}_{+}(\Gamma)$, then $G$ is both positively realizable and realizable for $P(n,k)$. Note however, that $G$ may be realizable and not be positively realizable. Also, by putting a distinct knot on each edge of $\Gamma$ we obtain an embedding $\Gamma^{\prime\prime}$ such that $\mathrm{TSG}(\Gamma^{\prime\prime})=\mathrm{TSG}_{+}(\Gamma^{\prime\prime})=\langle e\rangle$. Thus the trivial group is realizable and positively realizable for every $P(n,k)$. Hence we focus on non-trivial groups.
Since $P(n,k)$ is $3$-connected, we can use the following theorems.
@@
The above theorem allows us to assume realizable automorphisms are induced by finite order homeomorphisms of $S^{3}$, and then use the following simplified version of a theorem of P. A. Smith [15]. We let $\mathrm{fix}(h)$ denote the fixed point set of $h$.
Theorem 1.4 (Smith Theory).
Let $h$ be a finite order homeomorphism of $S^{3}$. If $h$ is orientation preserving then $\mathrm{fix}(h)$ is either the empty set or $S^{1}$, and if $h$ is orientation reversing then $\mathrm{fix}(h)$ is either two points or $S^{2}$.
We now list several other results that will be used. Note that $\mathrm{SO}(4)$ is the group of orientation preserving isometries of $S^{3}$ and $\mathrm{SO}(3)$ is the group of orientation preserving isometries of $S^{2}$. The Group Rigidity Theorem below follows from Proposition 3 of [9] together with the Geometrization Theorem [13].
@@
Let $\gamma$ be a $3$-connected graph and $G\leq\mathrm{TSG}_{+}(\Lambda)$ for some embedding $\Lambda$ of $\gamma$ in $S^{3}$. Then there is an embedding $\Gamma$ of $\gamma$ in $S^{3}$ such that $G\leq\mathrm{TSG}_{+}(\Gamma)$ and $\mathrm{TSG}_{+}(\Gamma)$ is induced by an isomorphic finite subgroup of $\mathrm{SO}(4)$.
@@
Let $G\leq SO(4)$ such that for every involution $g\in G$, we have $\mathrm{fix}(g)\cong S^{1}$ and no $h\in G$ with $g\neq h$ has $\mathrm{fix}(h)=\mathrm{fix}(g)$. Then $G$ is a subgroup of $D_{m}\times D_{m}$ for some odd $m$ or is a finite subgroup of $\mathrm{SO}(3)$.
@@
Then there is an embedding $\Gamma$ of the graph $\gamma$ in $S^{3}$ such that $\Gamma$ is setwise invariant under $G$.
We also use the website https://people.maths.bris.ac.uk/ matyd/GroupNames/ and the software Sage to identify the isomorphism classes of automorphism groups and their subgroups.
2. Realizability of $D_{n}$ and its subgroups for all $P(n,k)$
@@
We construct an embedding of $P(n,k)$ as follows. Let $0<r<R$. We begin with planar circles $C$ and $c$, centered at the origin, of radius $R$ and $r$, respectively. Let $u_{1}$ and $v_{1}$ be the points of intersection between a planar ray from the origin and $C$ and $c$, respectively. Next let $e_{1}$ denote the straight line segment between $u_{1}$ and $v_{1}$. Then $e_{1}$ lies on the ray. Let $f$ denote a rotation of $S^{3}$ of $\frac{2\pi}{n}$ about an axis through the origin which is perpendicular to the plane of the projection. Then for each $i$ let $u_{i}=f^{i}(u_{1})$, $v_{i}=f^{i}(v_{1})$, $e_{i}=f^{i}(e_{1})$. Thus we have embedded the spokes $\overline{u_{i}v_{i}}$ as the $e_{i}$. Next we embed the outer edges $\overline{u_{i}u_{i+1}}$ as straight line segments. We will embed the inner edges below.
Let $g$ denote a rotation of $\pi$ about a planar axis that contains the spoke $\overline{u_{n}v_{n}}$ if $n$ is odd and contains the two opposite spokes $\overline{u_{n}v_{n}}$ and $\overline{u_{\frac{n}{2}}v_{\frac{n}{2}}}$ if $n$ is even. Then the isometry group $D_{n}=\langle f,g\rangle$ leaves the $n$-gon of outer edges and the set of spokes invariant, inducing an isomorphic group on the embedded vertices of $P(n,k)$. Now we embed the inner edges $\overline{v_{i}v_{i+k}}$ so that the collection of edges $\{\overline{v_{i}v_{i+k}}|\medspace i\leq n\}$ is pairwise disjoint and setwise invariant under $D_{n}$. We can do this, for example, by making the crossings along each inner edge alternate, as illustrated for the embedding of $P(7,2)$ on the right of Figure 1. Let $\Gamma^{\prime}$ denote this embedding of $P(n,k)$. Thus $\Gamma^{\prime}$ is invariant under $\langle f,g\rangle$, and hence $D_{n}\leq\mathrm{TSG}_{+}(\Gamma^{\prime})$.
We obtain the embedding $\Gamma$ from $\Gamma^{\prime}$ by adding the invertible knot $4_{1}$ to each outer edge. Then $\Gamma$ is invariant under $\langle f,g\rangle$, and hence $D_{n}\leq\mathrm{TSG}_{+}(\Gamma)$. Let $L$ denote the $n$-gon of outer edges of $\Gamma^{\prime}$. Then $L$ is the only $n$-gon containing $n$ copies of the knot $4_{1}$, and hence any element of $\mathrm{TSG}_{+}(\Gamma)$ must take $L$ to itself. Since no non-trivial automorphism of $P(n,k)$ fixes every $u_{i}$, this implies that $D_{n}\leq\mathrm{TSG}_{+}(\Gamma)\leq\mathrm{TSG}_{+}(L)$. On the other hand, because the automorphism group of an $n$-gon is $D_{n}$, we have $\mathrm{TSG}_{+}(L)\leq D_{n}$, and hence $D_{n}=\mathrm{TSG}_{+}(\Gamma)$.
@@
Let $k^{2}\not\equiv\pm 1\pmod{n}$ where $(n,k)$ is non-exceptional. Then $\operatorname{Aut}(P(n,k))$ and all of its subgroups are positively realizable and hence realizable for $P(n,k)$.
@@ 3. The case $k^{2}\equiv 1\pmod{n}$
We use the following embedding $\Lambda$ of $P(n,k)$ in this and the next section. Let $U$ and $V$ denote the cores of complementary isometric solid tori in $S^{3}$. Hence $U$ and $V$ are geodesic circles, and for every point on one of these cores, its antipodal point is on the same core. Now, for every $u\in U$ and $v\in V$, there is a unique shortest geodesic $e$ joining $u$ and $v$, and the length of $e$ is less than $\pi$. Since $U$ and $V$ are geodesic circles, it follows that the interior of any such $e$ must be disjoint from $U\cup V$.
Suppose that $k^{2}\equiv\pm 1\pmod{n}$. Then $n$ and $k$ are relatively prime, and hence the inner edges of $P(n,k)$ form a single loop. The embedding $\Lambda$ is obtained as follows. We let $e_{1}$ be a geodesic of minimal length between $U$ and $V$, and let $u_{1}$ and $v_{1}$ be its endpoints on $U$ and $V$ respectively. Let $f$ be a glide rotation which rotates $U$ by $\frac{2\pi}{n}$ while rotating $V$ by $k\big{(}\frac{2\pi}{n}\big{)}$. Then for each $i$, define $e_{i+1}=f^{i}(e_{1})$, and define the endpoint of $e_{i+1}$ on $U$ to be $u_{i+1}$ and the endpoint on $V$ to be $v_{i+1}$. Thus $f(e_{i})=e_{i+1}$, $f(u_{i})=u_{i+1}$, and $f(v_{i})=v_{i+1}$.
Since $f$ rotates $U$ by $\frac{2\pi}{n}$, the points $u_{i}$ and $u_{i+1}$ are consecutive on $U$. Also, since $f$ rotates $V$ by $k\big{(}\frac{2\pi}{n}\big{)}$, we know that $f^{k}$ rotates $V$ by $k^{2}\big{(}\frac{2\pi}{n}\big{)}\equiv\frac{2\pi}{n}\pmod{n}$. Since $f^{k}(v_{i})=v_{i+k}$, this implies that $v_{i}$ and $v_{i+k}$ are consecutive on $V$. Thus we define the edges $\overline{u_{i}u_{i+1}}$ and $\overline{v_{i}v_{i+k}}$ to be minimal arcs on $U$ and $V$, respectively. It follows that the interiors of the $e_{i}$ are disjoint from $U$ and $V$. Furthermore, if we consider the open book decomposition of $S^{3}$ whose spine is $U$, then $f$ rotates the pages of this decomposition around $U$ and each $e_{i}$ is on a distinct page. Hence the $e_{i}$ are pairwise disjoint. This gives us an embedding $\Lambda$ of $P(n,k)$.
Now we orient $U$ and $V$ so that $\overrightarrow{u_{i}u_{i+1}}$ and $\overrightarrow{v_{i}v_{i+k}}$ are oriented positively.
@@
We start with the embedding $\Lambda$ of $P(n,k)$ described above and we add a $4_{1}$ knot to each edge of $U$ and $V$ to get an embedding $\Lambda^{\prime}$ of $P(n,k)$. Recall that $f$ is a glide rotation which rotates $U$ by $\frac{2\pi}{n}$ while rotating $V$ by $k\big{(}\frac{2\pi}{n}\big{)}$ both in the positive direction. Let $g$ be a rotation of $S^{3}$ by $\pi$ around an axis that contains $e_{1}=\overline{u_{1}v_{1}}$. Then $g$ takes $U$ and $V$ to themselves, reversing their orientations. Finally, let $h$ be an order $2$ rotation of $S^{3}$ interchanging the positively oriented $U$ with the positively oriented $V$, taking $u_{n}$ to $v_{n}$.
In order to show that $f$, $g$, and $h$ take $\Lambda^{\prime}$ to itself, we need to know that they take edges to edges. First, by definition we know that $f(e_{i})=e_{i+1}$. Also, for every $i$, we have $g(v_{1-ki})=v_{1+ki}$, and hence $g(v_{1-i})=g(v_{1-k^{2}i})=v_{1+k^{2}i}=v_{1+i}$. Since we also have $g(u_{1-i})=u_{1+i}$ and we know that the $e_{i}$ are minimal length geodesics, it follows that $g(e_{1-i})=e_{i+1}$. Finally, since $h$ interchanges $u_{n}$ and $v_{n}$ preserving orientation, we have $h(u_{i})=v_{ki}$, and hence $h(u_{ki})=v_{k^{2}i}=v_{i}$. Thus for every $i$, we have $h(e_{i})=e_{ki}$. It follows that $f$, $g$, and $h$ take $\Lambda^{\prime}$ to itself.
Now $f$ induces a rotation of order $n$ on $U$ and $V$, $g$ turns $U$ and $V$ over, and $h$ is an involution interchanging $U$ and $V$. Thus $\langle f,g,h\rangle$ induces $D_{n}\rtimes\mathbb{Z}_{2}$ on $\Lambda^{\prime}$, and hence $D_{n}\rtimes\mathbb{Z}_{2}\leq\mathrm{TSG}_{+}(\Lambda^{\prime})$. However, since every element of $\mathrm{TSG}_{+}(\Lambda)$ takes the set of edges $e_{i}$ to themselves, it follows that $\mathrm{TSG}_{+}(\Lambda^{\prime})\leq B(n,k)$. Now by [12], $B(n,k)=D_{n}\rtimes\mathbb{Z}_{2}$ and hence $\mathrm{TSG}_{+}(\Lambda^{\prime})=D_{n}\rtimes\mathbb{Z}_{2}$.
@@
If $(n,k)$ is non-exceptional then $\mathrm{Aut}(P(n,k))=B(n,k)$, implying the following.